how do you go about solving the following:
2x+y-z=5
3x-y+2z=-1
x-y-z=19
try to solve this system using either elimination or substitution:

Answer by Nate(3500)

3x - y + 2z = -1
-x + y + z = -19
2x + 3z = -20 from elimination
2x + y - z = 5
x - y - z = 19
3x - 2z = 24 from elimination
(2x + 3z = -20)2
(3x - 2z = 24)3
4x + 6z = -40
9x - 6z = 72
13x = 32
x = 32/13
plug:
3x - 2z = 24
3(32/13) - 2z = 24
96/13 - 2z = 312/13
-2z = 216/13
z = -108/13
plug:
x - y - z = 19
32/13 - y + 108/13 = 247/13
-y + 140/13 = 247/13
-y = 107/13
y = -107/13
z = -108/13
x = 32/13